A circle has it's center at (2,-1) and one end of the diameter at (5,1). Find the other end of the diameter..
that is one of the question i can't answer, please help me with that.. and please give me the formula for that.. thank you...

Answer by ewatrrr(24785)

 
Hi,
A circle has it's center at (2,-1) and one end of the diameter at (5,1).
(5,1)
(2,-1) m = 2/3 y=(2/3)x+ b -1=4/3+b b =-7/3 y = (2/3)x - 7/3 Green
Slope of the Line containing the Diameter is m = 2/3
y increases/decreases by 2, x increases/decreases by 3
other end of diameter is (-1,-3) down 2 and back 3 from center Pt(2,-1)